OpenStudy (anonymous):

Prove: For all integers n, if n^2 is odd, then n is odd. Use a proof by contraposition, as in Lemma 1.1. Let n be an integer. Suppose that n is even, i.e., n = blank for some integer k. Then n^2 = blank = 2( blank) is also even.

OpenStudy (anonymous):

Form the contrapositive of the given statement. That is, For all integers n, if n is not odd, then n2 is not odd But, we know that an integer is not odd if, and only if, it is even [by parity property]. So, the contrapositive becomes For all integer n, if n is even, then n^2 is even

OpenStudy (anonymous):

The prove the contrapositive using method of direct proof: Suppose n is an integer. We must show that n^2 is also even. By definition of even, we have n = 2k for some integer k.

OpenStudy (anonymous):

Then by substitution, we have n.n = (2k) . (2k) n.n= 4(k^2) n.n=2(2k^2) Because products of integers are integers and 2 and k are both integers.] Hence, we have a form: n . n = 2 . (some integer) and so by definition of even is n^2 is even.

OpenStudy (anonymous):

Therefore, the given statement is true by the logical equivalence between a statement and its contrapositve.

OpenStudy (anonymous):

I tried using 2 and the problem said that that was incorrect

OpenStudy (anonymous):

do you have to submit the answer electronically?

OpenStudy (anonymous):

yeah there is a box beside n= and then another after n to the second power and then another after =2( )

OpenStudy (anonymous):

can you screen shot it?

OpenStudy (anonymous):

Click on the attachment

OpenStudy (anonymous):

did you try n=2k for the first bit?

OpenStudy (anonymous):

then n=4k^2 = 2(2k^2)

OpenStudy (anonymous):

That worked chris00 I appreciate it

OpenStudy (anonymous):

: )